# Temporal Inductive Path Neural Network for Temporal Knowledge Graph Reasoning

### 08 Dec 2023 v2

@@ 4.4. Query-aware Temporal Path Processing
4.4.1. Temporal Path Aggregation Layer
We propose a temporal path aggregation layer for comprehensively aggregating the connected temporal edges between the query subject entity $s$ and all candidate object entities in $\mathcal{V}$ within the history temporal graph.
Note that a temporal path representation is specific to a particular query. Different queries can lead to the same path, but they have different representations (e.g. when two queries share the same subject entity but have different query relations). This insight design ensures each path representations are context-dependent and take into account the query’s unique characteristics, such as the query subject entity and query relation type, which can influence the meaning and relevance of the path in different contexts.
@@ 4.4.1. Temporal Path Aggregation Layer
A temporal path is formed by logically aggregating multiple real existing paths from the history temporal graph. Each path consists of multiple temporal edges. Although the temporal edge exists in the form of relation-timestamp, it is still based on a specific relationship $r\in\mathcal{R}$. Therefore, we initialize a trainable representation for the relation set, which serves as a foundational feature when processing temporal edge features during temporal path aggregation with query awareness.
Here, we introduce and discuss the learnable relation representation and initialization of temporal path representation.
@@
    (4) $\textbf{h}^{0}_{o}\leftarrow\begin{cases}\textbf{r}&if\ o\Leftrightarrow s,\\ \vec{0}&if\ o\nLeftrightarrow s.\end{cases}$
@@
(5) $\textbf{h}_{o}^{l}=\textsc{Agg}\bigg{(}\Big{\{}\textsc{Tmsg}\big{(}\textbf{h}_{z}^{l-1},\textbf{w}_{r}(z,p_{\tau},o)\big{)}\Big{|}(z,p_{\tau},o)\in\hat{G}_{\prec t+1}\Big{\}}\bigg{)},$
@@
4.4.2. Aggregation Operating
Equation 5 describes the iterative aggregation process of $\textbf{h}_{o}$ in the $l$-th layer. Here we provide explanations for the temporal edge merging function $\textsc{Tmsg}(\cdot)$ and path aggregation function $\textsc{Agg}(\cdot)$, respectively.
Intuitively, the $\textsc{Tmsg}(\cdot)$ function is responsible for iteratively searching for paths with a candidate object entity $o$ as the endpoint, continuously propagating information along the temporal edges. On the other hand, the $\textsc{Agg}(\cdot)$ function is responsible for aggregating the information obtained from each iteration along the temporal edges. Due to the initial state of $\textbf{H}_{o}^{0}$, where only $\textbf{h}_{s}^{0}$ is given the initial information, after $\omega$ iterations of aggregation, $\textbf{h}_{o}^{\omega}$ captures the comprehensive semantic connection and temporal dependency information of multiple paths from subject node $s$ to object candidate $o$ in the history temporal graph $\hat{G}_{\prec t+1}$.
@@ 4.4.2. Aggregation Operating
    It takes the current temporal path representation $\textbf{h}_{z}^{l-1}$, query relation representation r, and the temporal edge information as input and calculates the updated potential path information for the candidate object $o$. Similar to message passing in knowledge graphs, we have adopted a vectorized multiplication method (Yang et al., 2014) to achieve feature propagation in history temporal graph, following
    (6) $\textsc{Tmsg}\big{(}\textbf{h}_{z}^{l-1},\textbf{w}_{r}(z,p_{\tau},o)\big{)}=\textbf{h}_{z}^{l-1}\otimes\textbf{w}_{r}(z,p_{\tau},o),$
@@
    where the operator $\otimes$ is defined as element-wise multiplication between $\textbf{h}_{z}^{l-1}$ and $\textbf{w}_{r}(z,p_{\tau},o)$. The vectorized multiplication can be understood as scaling $\textbf{h}_{z}^{l-1}$ by $\textbf{w}_{r}(z,p_{\tau},o)$ in our temporal edges merging (Yang et al., 2014). As mentioned earlier, $\textbf{w}_{r}(z,p_{\tau},o)$ represents the query-aware temporal representation of a temporal edge $(z,p_{\tau},o)$, which needs to ensure its relevance to the query, enabling effective capturing of relation features and temporal characteristics. Therefore, the approach of temporal relation encoding is proposed for the temporal edge representation.
@@
    Temporal Relation Encoding. Temporal edge representation should align with the query’s characteristics, facilitating more contextually relevant and informative feature propagation. Specifically, we take into account the temporal edge features from the query relation representation, semantic and temporal characteristics of the temporal edges in the history temporal graph, generating comprehensive embedding with both static and temporal characteristics. Formally, $\textbf{w}_{r}(z,p_{\tau},o)$ can be derived as
    (7) $\textbf{w}_{r}(z,p_{\tau},o)=g\Bigg{(}\bm{\Psi}_{r}(p)\bigg{|}\bigg{|}\bm{\Upsilon}(\Delta\tau)\Bigg{)},$
@@
    where $\bm{\Psi}_{r}(p)\in\mathbb{R}^{d}$ denotes the query relation $r$-aware basic static representation of edge type $p$, $\bm{\Upsilon}(\Delta\tau)\in\mathbb{R}^{d}$ denotes the temporal embedding of the temporal edge type $p_{\tau}$, ||||| | denotes the operator of embedding concatenation, and $g(\cdot)$ is a feed-forward neural network. For the basic static representation of $p$, we obtain it through a linear transformation. Based on the query relation representation $\mathbf{r}$, we map a representation as the static semantic information of the temporal edge type $p$, following
    (8) $\bm{\Psi}_{r}(p)=\textbf{W}_{p}\textbf{r}+\textbf{b}_{p},$
@@
    where $\textbf{W}_{p}$ and $\textbf{b}_{p}$ are learnable parameters and serve as the weights and biases of the linear transformation, respectively, which makes the basic static information of the temporal edges derived from the given query relation embedding, ensuring the awareness of the query. And for the temporal embedding of the temporal edge, we use generic time encoding (Xu et al., 2020b) to model the temporal information in temporal edges, following
    (9) $\Delta\tau=|\tau-t_{q}|,$
    (10) $\bm{\Upsilon}(\Delta\tau)=\sqrt{\frac{1}{d}}\ \ \Big{[}cos(\bm{w}_{1}\Delta\tau+\bm{\phi}_{1}),cos(\bm{w}_{2}\Delta\tau+\bm{\phi}_{2}),\cdots,cos(\bm{w}_{d}\Delta\tau+\bm{\phi}_{d})\ \Big{]},$
@@
    It considers the accumulated information from the previous iteration and the newly propagated information along the temporal edge to update the temporal path representation for a candidate object. We adopt principal neighborhood aggregation (PNA) proposed in (Corso et al., 2020), which leverages multiple aggregators (namely mean, maximum, minimum, and standard deviation) to learn joint feature, since previous work has verified its effectiveness (Zhu et al., 2022)(Miao et al., 2022). We also consider the traditional aggregation function as a comparison, such as sum, mean, and max, which will be introduced in detail in Section 5.3.
Finally, after the $\omega$-th iteration of aggregation, $\textbf{H}_{\mathcal{V}}^{\omega}$ will obtain the representation of the temporal paths from query subject $s$ to all candidate objects in $\mathcal{V}$, with comprehensive semantic connection and temporal dependency information of multiple paths.
4.5. Learning and Inference
TiPNN models the query-aware temporal path feature by aggregation process within the history temporal graph. By capturing comprehensive embedding with both static and temporal characteristics of temporal edges, and aggregating multiple paths that consist of temporal edges, TiPNN learns a comprehensive representation of temporal path. Different from most previous models, we utilize the temporal edge features in a path from query subject to candidate object, without considering any node embedding during modeling processing, thus TiPNN can solve the inductive setting, which will be presented in Section 5.2 in detail.
4.5.1. Score Function
Here we show how to apply the final learned temporal paths representation to the temporal knowledge graph reasoning.
Given query subject $s$ and query relation $r$ at timestamp $t+1$, after obtaining temporal paths representation $\textbf{H}_{(s,r)\rightarrow{\mathcal{V}}}^{\omega}$, we predict the conditional likelihood of the future object candidate $o\in\mathcal{V}$ using $\textbf{h}_{(s,r)\rightarrow{o}}^{\omega}\in\textbf{H}_{(s,r)\rightarrow{\mathcal{V}}}^{\omega}$, following:
(11) $p(o|s,r)=\sigma\Bigg{(}\mathcal{F}\Big{(}\textbf{h}_{(s,r)\rightarrow{o}}^{\omega}\ \big{|}\big{|}\ \bm{r}\Big{)}\Bigg{)},$
@@ 4.5.1. Score Function
(12) $p(s|o,r^{-1})=\sigma\Bigg{(}\mathcal{F}\Big{(}\textbf{h}_{(o,r^{-1})\rightarrow{s}}^{\omega}\ \big{|}\big{|}\ \bm{r}^{-1}\Big{)}\Bigg{)}.$
4.5.2. Parameter Learning
Reasoning on a given query can be seen as a binary classification problem. The objective is to minimize the negative log-likelihood of positive and negative triplets, as shown in Equation 13.
In the process of generating negative samples, we follow the Partial Completeness Assumption  (Galárraga et al., 2013). Accordingly, for each positive triplet in the reasoning future triplet set, we create a corresponding negative triplet by randomly replacing one of the entities with a different entity. It ensures that the negative samples are derived from the future triplet set at the same timestamp, allowing the model to effectively learn to discriminate between correct and incorrect predictions at the same future timestamp.
(13) $\mathcal{L}_{TKG}=-\log p(s,r,o)-\sum_{j=1}^{n}\frac{1}{n}log(1-p(\overline{s_{j}},r,\overline{o_{j}})),$
@@ 4.5.2. Parameter Learning
(14) $\mathcal{L}_{REG}=\Big{\|}\ \textbf{R}^{T}\textbf{R}-\alpha\textbf{I}\ \Big{\|}$
@@
(15) $\mathcal{L}=\mathcal{L}_{TKG}+\mathcal{L}_{REG}.$
4.6. Complexity Analysis
The score of each temporal path from the query subject and candidate object can be calculated in parallel operation. Therefore, to see the complexity of the proposed TiPNN, we analyze the computational complexity of each query. The major operation in TiPNN is query-aware temporal path processing, and each aggregation iteration contains two steps: temporal edge merging function $\textsc{Tmsg}(\cdot)$ and path aggregation function $\textsc{Agg}(\cdot)$ (as shown in Equation 5). We use $|\mathcal{E}|$ to denote the maximum number of concurrent facts in the historical subgraph sequence, and $|\mathcal{V}|$ to denote the cardinality of the entity set. $\textsc{Tmsg}(\cdot)$ has the time complexity of $O(m|\mathcal{E}|)$ since it performs Equation 6 for all edges in the history temporal graph, where $m$ denotes the length of the history used for reasoning. $\textsc{Agg}(\cdot)$ has the time complexity of $O(|\mathcal{V}|)$ since the aggregation method we adopt will perform for every node. And $\textsc{Agg}(\cdot)$ is executed after $\textsc{Msg}(\cdot)$ at each aggregation iteration, so the time complexity is finally $O(\omega(m|\mathcal{E}|+|\mathcal{V}|))$ when the number of aggregation layers is $\omega$.
5. Experiments
In this section, we carry out a series of experiments to assess the effectiveness and performance of our proposed TiPNN. Through these experiments, we aim to address and answer the following research questions:
@@ 5. Experiments
5.1. Experimental Settings
5.1.1. Datasets
We conduct extensive experiments on four representative temporal knowledge graph datasets, namely, ICEWS18 (Jin et al., 2020), GDELT (Leetaru and Schrodt, 2013), WIKI (Leblay and Chekol, 2018), and YAGO (Mahdisoltani et al., 2014). These datasets are widely used in the field of temporal knowledge reasoning due to their diverse temporal information. The ICEWS18 dataset is collected from the Integrated Crisis Early Warning System (Boschee et al., 2015), which records various events and activities related to global conflicts and crises, offering valuable insights into temporal patterns of such events. The GDELT dataset (Jin et al., 2019) is derived from the Global Database of Events, Language, and Tone, encompassing a broad range of events across the world and enabling a comprehensive analysis of global events over time. WIKI (Leblay and Chekol, 2018) and YAGO (Mahdisoltani et al., 2014) are two prominent knowledge bases containing factual information with explicit temporal details. WIKI covers a wide array of real-world knowledge with timestamps, while YAGO offers a structured knowledge base with rich semantic relationships and temporal annotations. And we focus on the subsets of WIKI and YAGO that have yearly granularity in our experiments. The details of the datasets are provided in Table 2. To ensure fair evaluation, we adopt the dataset splitting strategy employed in  (Jin et al., 2020). The dataset is partitioned into three sets: training, validation, and test sets, based on timestamps. Specifically, (timestamps of the train) $<$ (timestamps of the valid) $<$ (timestamps of the test).
Table 2. Statistics of Datasets ($\mathcal{E}_{train}$, $\mathcal{E}_{valid}$, $\mathcal{E}_{test}$ are the numbers of facts in training, validation, and test sets.).

### 25 Jan 2024 v3

@@ 3.4 Query-aware Temporal Path Processing
3.4.1 Temporal Path Aggregation Layer
We propose a temporal path aggregation layer for comprehensively aggregating the connected temporal edges between the query subject entity $s$ and all candidate object entities in $\mathcal{V}$ within the history temporal graph. Note that a temporal path representation is specific to a particular query. Different queries can lead to the same path, but they have different representations (e.g. when two queries share the same subject entity but have different query relations). This design ensures each path representations are context-dependent and take into account the query’s unique characteristics, such as the query subject entity and query relation type, which can influence the meaning and relevance of the path in different contexts.
@@ 3.4.1 Temporal Path Aggregation Layer
A temporal path is formed by logically aggregating multiple real existing paths from the history temporal graph. Each path consists of multiple temporal edges. Although the temporal edge exists in the form of relation-timestamp, it is still based on a specific relationship $r\in\mathcal{R}$. Therefore, we initialize a trainable representation for the relation set, which serves as a foundational feature when processing temporal edge features during temporal path aggregation with query awareness. Here, we introduce and discuss the learnable relation representation and initialization of temporal path representation.
  • 1.
@@
    $\textbf{h}^{0}_{o}\leftarrow\begin{cases}\textbf{r}&if\ o\Leftrightarrow s,\\ \vec{0}&if\ o\nLeftrightarrow s.\end{cases}$ (4)
@@
$\textbf{h}_{o}^{l}=\textsc{Agg}\bigg{(}\Big{\{}\textsc{Tmsg}\big{(}\textbf{h}_{z}^{l-1},\textbf{w}_{r}(z,p_{\tau},o)\big{)}\Big{|}(z,p_{\tau},o)\in\hat{G}_{\prec t+1}\Big{\}}\bigg{)},$ (5)
@@
However, to emphasize the distinction from DaeMon dong2023adaptive , it is important to note that in DaeMon, the acquisition of temporal paths is learned independently on each isolated historical subgraph (i.e., $G_{i}$, where $i<t+1$). Consequently, the final temporal path representation requires stacking $m$-times (i.e., history length) updates of $\textbf{h}_{o}$ in DaeMon.
Figure 3: An illustrative diagram of the proposed TiPNN model for query-aware temporal path processing. For a given query $(s,r,?)$ at future timestamp, TiPNN engages in temporal path processing within the constructed history temporal graph $\hat{G}$ to perform prediction for the future timestamp. The temporal path feature $\textbf{H}_{\mathcal{V}}$ is iteratively learned and updated by the query-aware $\omega$-layers aggregation neural network. In each layer, the temporal edges in $\hat{G}$, enriched with temporal information, are individually modeled for basic static representation and temporal representation through $\bm{\Psi}_{r}$ and $\bm{\Upsilon}$, respectively. Subsequently, temporal edges are merged using $\textsc{Tmsg}(\cdot)$, followed by aggregation through $\textsc{Agg}(\cdot)$ to obtain the feature of the current layer’s temporal path.
3.4.2 Aggregation Operating
Equation 5 describes the iterative aggregation process of $\textbf{h}_{o}$ in the $l$-th layer. Here we provide explanations for the temporal edge merging function $\textsc{Tmsg}(\cdot)$ and path aggregation function $\textsc{Agg}(\cdot)$, respectively. Intuitively, the $\textsc{Tmsg}(\cdot)$ function is responsible for iteratively searching for paths with a candidate object entity $o$ as the endpoint, continuously propagating information along the temporal edges. On the other hand, the $\textsc{Agg}(\cdot)$ function is responsible for aggregating the information obtained from each iteration along the temporal edges. Due to the initial state of $\textbf{H}_{o}^{0}$, where only $\textbf{h}_{s}^{0}$ is given the initial information, after $\omega$ iterations of aggregation, $\textbf{h}_{o}^{\omega}$ captures the comprehensive semantic connection and temporal dependency information of multiple paths from subject $s$ to object candidate $o$ in the history temporal graph $\hat{G}_{\prec t+1}$.
  • 1.
@@ 3.4.2 Aggregation Operating
    It takes the current temporal path representation $\textbf{h}_{z}^{l-1}$, query relation representation r, and the temporal edge information as input and calculates the updated potential path information for the candidate object $o$. Similar to message passing in knowledge graphs, we have adopted a vectorized multiplication method yang2014embedding to achieve feature propagation in history temporal graph, following
    $\textsc{Tmsg}\big{(}\textbf{h}_{z}^{l-1},\textbf{w}_{r}(z,p_{\tau},o)\big{)}=\textbf{h}_{z}^{l-1}\otimes\textbf{w}_{r}(z,p_{\tau},o),$ (6)
@@
    where the operator $\otimes$ is defined as element-wise multiplication between $\textbf{h}_{z}^{l-1}$ and $\textbf{w}_{r}(z,p_{\tau},o)$. The vectorized multiplication can be understood as scaling $\textbf{h}_{z}^{l-1}$ by $\textbf{w}_{r}(z,p_{\tau},o)$ in our temporal edges merging yang2014embedding . $\textbf{w}_{r}(z,p_{\tau},o)$ represents the query-aware temporal representation of a temporal edge $(z,p_{\tau},o)$, which needs to ensure its relevance to the query. Different from DaeMon dong2023adaptive , in the history temporal graph, each edge is appended with temporal information, preventing the consideration of features solely based on static relation types. Therefore, the approach of temporal relation encoding is proposed for the temporal edge representation.
@@
    Temporal Relation Encoding. As mentioned earlier, temporal edge representation should align with the query’s characteristics. We take into account the temporal edge features from the query relation representation, semantic and temporal characteristics of the temporal edges in the history temporal graph, generating comprehensive embedding with both static and temporal characteristics. Formally, $\textbf{w}_{r}(z,p_{\tau},o)$ can be derived as
    $\textbf{w}_{r}(z,p_{\tau},o)=g\Bigg{(}\bm{\Psi}_{r}(p)\bigg{|}\bigg{|}\bm{\Upsilon}(\Delta\tau)\Bigg{)},$ (7)
@@
    where $\bm{\Psi}_{r}(p)\in\mathbb{R}^{d}$ denotes the query relation $r$-aware basic static representation of edge type $p$, $\bm{\Upsilon}(\Delta\tau)\in\mathbb{R}^{d}$ denotes the temporal embedding of the temporal edge type $p_{\tau}$, ||||| | denotes the operator of concatenation, and $g(\cdot)$ is a feed-forward neural network. For the basic static representation of $p$, following dong2023adaptive , we obtain it through a linear transformation. Based on the query relation representation $\mathbf{r}$, we map a representation as the static semantic information of the temporal edge type $p$, following
    $\bm{\Psi}_{r}(p)=\textbf{W}_{p}\textbf{r}+\textbf{b}_{p},$ (8)
@@
    where $\textbf{W}_{p}$ and $\textbf{b}_{p}$ are learnable parameters and serve as the weights and biases of the linear transformation, respectively, which makes the basic static information of the temporal edges derived from the given query relation embedding, ensuring the awareness of the query. And for the temporal embedding of the temporal edge, we use generic time encoding xu2020inductive to model the temporal information in temporal edges, following
    $\Delta\tau=|\tau-t_{q}|,$ (9)
    $\bm{\Upsilon}(\Delta\tau)=\sqrt{\frac{1}{d}}\ \ \Big{[}cos(\bm{w}_{1}\Delta\tau+\bm{\phi}_{1}),cos(\bm{w}_{2}\Delta\tau+\bm{\phi}_{2}),\cdots,cos(\bm{w}_{d}\Delta\tau+\bm{\phi}_{d})\ \Big{]},$ (10)
@@
    It considers the accumulated information from the previous iteration and the newly propagated information along the temporal edge to update the temporal path representation for a candidate object. We adopt principal neighborhood aggregation (PNA) proposed in corso2020principal , which leverages multiple aggregators (namely mean, maximum, minimum, and standard deviation) to learn joint feature, since previous work has verified its effectiveness zhu2022neural ; pmlr-v162-miao22a . We also consider the traditional aggregation function as a comparison, such as sum, mean, and max, which will be introduced in Section 4.3.
Finally, after the $\omega$-th iteration of aggregation, $\textbf{H}_{\mathcal{V}}^{\omega}$ will obtain the representation of the temporal paths from query subject $s$ to all candidate objects in $\mathcal{V}$. Here we should note that, unlike the previous work dong2023adaptive , TiPNN integrates all historical subgraphs into a unified history temporal graph $\hat{G}$. When handling temporal path features, any temporal edge $(z,p_{\tau},o)\in\hat{G}$ is associated with timestamp information. Additionally, the static representation of $p_{\tau}$ and its temporal representation are respectively processed by $\bm{\Psi}$ and $\bm{\Upsilon}$. Consequently, TiPNN eliminates the need, as in DaeMon, for a separate consideration of cross-temporal path fusion. The resulting $\textbf{H}_{\mathcal{V}}^{\omega}$ already encompasses comprehensive semantic connections and temporal dependency information from multiple paths in history.
3.5 Learning and Inference
TiPNN models the query-aware temporal path feature by aggregation process within the history temporal graph. By capturing comprehensive embedding with both static and temporal characteristics of temporal edges, and aggregating multiple paths that consist of temporal edges, TiPNN learns a comprehensive representation of temporal path. Different from most previous models, we utilize the temporal edge features in a path from query subject to candidate object, without considering any entity embedding during modeling processing, thus TiPNN can solve the inductive setting, which will be presented in Section 4.2 in detail.
3.5.1 Score Function
Here we show how to apply the final learned temporal paths representation to the temporal knowledge graph reasoning. Given query subject $s$ and query relation $r$ at timestamp $t+1$, after obtaining temporal paths representation $\textbf{H}_{(s,r)\rightarrow{\mathcal{V}}}^{\omega}$, we predict the conditional likelihood of the future object candidate $o\in\mathcal{V}$ using $\textbf{h}_{(s,r)\rightarrow{o}}^{\omega}\in\textbf{H}_{(s,r)\rightarrow{\mathcal{V}}}^{\omega}$, following:
$p(o|s,r)=\sigma\Bigg{(}\mathcal{F}\Big{(}\textbf{h}_{(s,r)\rightarrow{o}}^{\omega}\ \big{|}\big{|}\ \bm{r}\Big{)}\Bigg{)},$ (11)
@@ 3.5.1 Score Function
$p(s|o,r^{-1})=\sigma\Bigg{(}\mathcal{F}\Big{(}\textbf{h}_{(o,r^{-1})\rightarrow{s}}^{\omega}\ \big{|}\big{|}\ \bm{r}^{-1}\Big{)}\Bigg{)}.$ (12)
3.5.2 Parameter Learning
Reasoning on a given query can be seen as a binary classification problem. The objective is to minimize the negative log-likelihood of positive and negative triplets, as shown in Equation 13. In the process of generating negative samples, we follow the Partial Completeness Assumption  galarraga2013amie . Accordingly, for each positive triplet in the reasoning future triplet set, we create a corresponding negative triplet by randomly replacing one of the entities with a different entity. It ensures that the negative samples are derived from the future triplet set at the same timestamp, allowing the model to effectively learn to discriminate between correct and incorrect predictions at the same future timestamp.
$\mathcal{L}_{TKG}=-\log p(s,r,o)-\sum_{j=1}^{n}\frac{1}{n}log(1-p(\overline{s_{j}},r,\overline{o_{j}})),$ (13)
@@ 3.5.2 Parameter Learning
$\mathcal{L}_{REG}=\Big{\|}\ \textbf{R}^{T}\textbf{R}-\alpha\textbf{I}\ \Big{\|}$ (14)
@@
$\mathcal{L}=\mathcal{L}_{TKG}+\mathcal{L}_{REG}.$ (15)
3.6 Complexity Analysis
The score of each temporal path from the query subject and candidate object can be calculated in parallel operation. Therefore, to see the complexity of the proposed TiPNN, we analyze the computational complexity of each query. The major operation in TiPNN is query-aware temporal path processing, and each aggregation iteration contains two steps: temporal edge merging function $\textsc{Tmsg}(\cdot)$ and path aggregation function $\textsc{Agg}(\cdot)$ (as shown in Equation 5). We use $|\mathcal{E}|$ to denote the maximum number of concurrent facts in the historical subgraph sequence, and $|\mathcal{V}|$ to denote the cardinality of the entity set. $\textsc{Tmsg}(\cdot)$ has the time complexity of $O(m|\mathcal{E}|)$ since it performs Equation 6 for all edges in the history temporal graph, where $m$ denotes the length of the history used for reasoning. $\textsc{Agg}(\cdot)$ has the time complexity of $O(|\mathcal{V}|)$ since the aggregation method we adopt will perform for every entity. And $\textsc{Agg}(\cdot)$ is executed after $\textsc{Tmsg}(\cdot)$ at each aggregation iteration, so the time complexity is finally $O(\omega(m|\mathcal{E}|+|\mathcal{V}|))$ when the number of aggregation layers is $\omega$.
4 Experiments
In this section, we carry out a series of experiments to assess the effectiveness and performance of proposed TiPNN. Through these experiments, we aim to address and answer the following research questions:
  • 1.
@@ 4 Experiments
4.1 Experimental Settings
4.1.1 Datasets
We conduct extensive experiments on four representative temporal knowledge graph datasets, namely, ICEWS18 jin2020Renet , GDELT 2013GDELT , WIKI 2018Deriving , and YAGO 2013YAGO3 . These datasets are widely used in the field of temporal knowledge reasoning due to their diverse temporal information. The ICEWS18 dataset is collected from the Integrated Crisis Early Warning System DVN/28075_2015 , which records various events and activities related to global conflicts and crises, offering valuable insights into temporal patterns of such events. The GDELT dataset jin2019recurrent is derived from the Global Database of Events, Language, and Tone, encompassing a broad range of events across the world and enabling a comprehensive analysis of global events over time. WIKI 2018Deriving and YAGO 2013YAGO3 are two prominent knowledge bases containing factual information with explicit temporal details. WIKI covers a wide array of real-world knowledge with timestamps, while YAGO offers a structured knowledge base with rich semantic relationships and temporal annotations. And we focus on the subsets of WIKI and YAGO that have yearly granularity in our experiments. The details of the datasets are provided in Table 2. To ensure fair evaluation, we adopt the dataset splitting strategy employed in  jin2020Renet . The dataset is partitioned into three sets: training, validation, and test sets, based on timestamps. Specifically, (timestamps of the train) $<$ (timestamps of the valid) $<$ (timestamps of the test).
Table 2: Statistics of Datasets ($\mathcal{E}_{train}$, $\mathcal{E}_{valid}$, $\mathcal{E}_{test}$ are the numbers of facts in training, validation, and test sets.).